OpenStudy (anonymous):

x ^{sqrt{logx}}=10^{8}

OpenStudy (mertsj):

\[x ^{\sqrt{\log_{10}x }}=10^8\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Can you lead me through the problem?

OpenStudy (mertsj):

Well, by inspection, we can figure out that x is 10^4. Can you see that?

OpenStudy (anonymous):

Yes, I understand that but I do not know what to do after?

OpenStudy (mertsj):

\[\sqrt{\log_{10}x }\log_{10}x=8 \]

OpenStudy (anonymous):

So log10 x = 8?

OpenStudy (mertsj):

Then if we square both sides we have: \[\log_{10}x(\log_{10}x)^2=64 \]

OpenStudy (mertsj):

\[(\log_{10}x )^3=64\]

OpenStudy (mertsj):

Now take the third root of both sides:

OpenStudy (mertsj):

\[\log_{10}x=4 \]

OpenStudy (mertsj):

Now put it into its exponential form

OpenStudy (anonymous):

Thank you so much!

OpenStudy (mertsj):

yw